# All-Sky Kinematics of the Distant Halo: The Reflex Response to the LMC

### v1

@@ 4.4 Local Effects of the LMC
To more closely examine the local velocity perturbations due to the LMC, we compare selected observed velocity trends to the simulations of GC19 in Figure 12. The observed datapoints in Figure 12 are directly measured from our stellar dataset, i.e., they are independent of the $v_{\mathrm{travel}}$ model described above. The top panel shows the mean Galactocentric radial velocity along the $L_{\mathrm{MS}}$ coordinate, which is roughly aligned with the past orbit of the LMC (Nidever et al., 2008). The extent of this plot only contains stars in the southern Galactic hemisphere. The overall trend of decreasing $v_{\mathrm{GSR}}$ versus $L_{\mathrm{MS}}$ is a signature of the global reflex response (red line), but there are higher-order ‘wiggles’ in the trend from N-body MW+LMC simulations (blue lines) that signal local effects from the LMC. It is clear that even along the relatively precisely-measured radial velocity component, a much larger sample of stars will be required to disentangle the kinematic signature of the local wake from the global reflex motion.
@@ 5.1 Context and Limitations
Before discussing the conclusions of our work, it is important to highlight the limitations of our dataset and methodology. As Figure 2 illustrates, our MagE survey has expanded spectroscopic coverage of the outer halo to the southern hemisphere, a major improvement over past samples. However, the bottom panel of Figure 2 also shows that the MagE survey is much more efficient at identifying stars beyond $\gtrsim 60$ kpc than corresponding northern surveys (H3 and SEGUE). Therefore, there is a strong need to observe more $\gtrsim 60$ kpc stars in the northern hemisphere, to achieve homogeneous sky coverage at these distances. Regardless, we have demonstrated (see Figure 4 and Appendix A) that the deeper southern coverage should not negatively affect the main measurements in our paper.
@@
The $v_{\mathrm{travel}}$ measured here differs in the sign of $v_{\mathrm{Y}}$ compared to the past measurements of PP21 and YPP24, although Figure 8 illustrates that the discrepancy is not drastic relative to the measurement uncertainties. As discussed earlier, one plausible culprit is the limited distance and sky coverage of the stellar samples used by PP21, YPP24. As the appendices in YPP24 demonstrate, this can influence the inferred direction of the reflex velocity. Our all-sky radial velocity map (middle panel of Figure 3) illustrates that there is a strong dipole signal oriented towards the LMC, and it is apparent that a lack of southern data would make the dipole direction more ambiguous.
@@
That being said, our agreement with simulations is by no means perfect, with the amplitude of our measured $v_{\mathrm{Y}}$ being larger than the simulations of GC19 and the Sgr measurements of Vasiliev et al. (2021). One caveat is that our inclusion of southern data could make our measurements more susceptible to local halo perturbations of the LMC. Although the simulations of GC19 suggest that the all-sky velocity structure is quite smooth and dominated by the reflex motion itself, any deviations from the simulation (or unresolved substructure from the LMC halo itself) could bias our measurement.
@@
Finally, an implicit assumption in our measurements — and the comparison to simulations — is that there is a relatively smooth component of the MW halo, relative to which a reflex velocity can be measured. We have carefully removed known substructures including dwarf galaxies and the Sagittarius Stream from our dataset. However, additional unknown substructures may be present in the data. When fitting our $v_{\mathrm{travel}}$ model, we iteratively clip stars that disagree with the model by $>3\sigma$. These cuts should prevent small-scale undetected substructures from biasing our overall measurement. Regardless, given the emerging picture of the various constituents of the outer halo, it would be desirable for future simulations of the MW+LMC interaction to incorporate more realistic and structured models of the MW halo. In addition to including the Sagittarius Stream (e.g., Vasiliev et al., 2021), simulations could utilize the latest observations of the Galactic halo’s spatially-varying density profile (e.g., Han et al., 2022; Amarante et al., 2024) and velocity ellipsoid (e.g., Bird et al., 2021, Figures 10-11 in this work).
@@ 5.2 The Dynamical Disequilibrium of the Outer Halo
Prior to any detailed modelling, §$\S$3 and Figures 3-5 in this work clearly illustrate the dramatic departure from any notion of ‘dynamical equilibrium’ beyond $\approx 50$ kpc. In §$\S$4 we use a physical model to demonstrate that the kinematic disequilibrium of the Galactic outskirts can be explained by the gravitational influence of the LMC. The kinematic influence of the LMC on our sample of stars is notably strong, suggesting an LMC dark matter halo on the massive side of recent estimates, $\gtrsim 1.8\times 10^{11}\,M_{\odot}$.
@@ 5.3 The Elusive Dark Matter Wake of the LMC
The dynamical friction wake of the LMC is predicted to leave imprints in both the density and kinematics of halo stars along the past orbit of the LMC. Although this wake has been detected in the density of stars — forming the so-called ‘Pisces Plume’ overdensity — a confident detection of the wake’s kinematic signature remains elusive (Belokurov et al., 2019; Conroy et al., 2021). In Figure 12 we search for the LMC’s local kinematic perturbation in radial and tangential velocities, over and above the bulk reflex motion. However, even in the relatively well-measured radial velocity direction, the detection is weak to non-existent given current observational errors.
@@ 6 Conclusions
We have assembled the largest all-sky dataset of $850$ luminous red giant stars from $40-160$ kpc with full 6D kinematics, including $\approx 300$ new stars in the southern hemisphere from our own tailored spectroscopic survey. In this work, we utilize this dataset to measure the motion of the inner galaxy relative to the outer halo, and find that it is induced by a massive and recently-infalling LMC. Our main conclusions are summarized below.
  • We construct an all-sky dataset of $850$ red giant stars beyond $40$ kpc using spectroscopy from H3, SEGUE, and our own MagE outer halo survey. All spectra are homogeneously analyzed with the same stellar parameter pipeline, resulting in the largest all-sky catalog of distant halo stars with 6D kinematic information (Figures 1-2).
@@
  • We fit the 6D positions and velocities of 800 stars from $12040<r_{\mathrm{gal}}<$ 120 to infer the amplitude and direction of the inner MW’s motion relative to distant halo stars. The apex direction of this travel velocity points towards the past orbit of the LMC, close to its present-day location (Figure 6, top). The amplitude of the travel velocity grows from $\sim 0$ at $r_{\mathrm{gal}}=40$ up to $50\pm 20$ km s-1 at $r_{\mathrm{gal}}=120$ kpc (Figure 6, bottom). A 3-D schematic of the measurement is shown in Figure 9 (animated version here).
  • The $v_{\mathrm{travel}}$ direction and amplitude measured in this work remarkably align with predictions from N-body simulations of the MW halo’s response to the LMC. The measured amplitude of the reflex velocity is most consistent with LMC halo masses above $\gtrsim 1.8\times 10^{11}\,M_{\odot}$, based on the simulations of GC19 (Figure 6).
@@
  • The intrinsic velocity dispersion of the MW halo declines as a function of distance in along all three spherical coordinates. After accounting for the LMC-induced reflex motion, outer halo stars exhibit significantly radially-biased motions, with a near-constant profile of the velocity anisotropy parameter $\beta\approx 0.6$ out to $120$ kpc (Figure 11).
@@
VC gratefully acknowledges a Peirce Fellowship from Harvard University. RPN acknowledges support for this work provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. It is a pleasure to thank Nitya Kallivayalil, David Nidever, Theo O’Neill, Ekta Patel, Michael Petersen, Yanjun Sheng, Eugene Vasiliev, and Rashid Yaaqib for insightful conversations and feedback. This work benefited from discussions at the 2024 “Milky Clouds over Manhattan” Workshop at the Flatiron Institute Center for Computational Astrophysics. We are indebted to the CfA and MIT telescope time allocation committees for enabling our long-term survey of the Galactic outskirts. We thank the staff at Las Campanas Observatory — including Yuri Beletsky, Carla Fuentes, Jorge Araya, Hugo Rivera, Alberto Pastén, Mauricio Martínez, Roger Leiton, Matías Díaz, Carlos Contreras, and Gabriel Prieto — for their invaluable assistance. The H3 Survey is funded in part by NSF grant NSF AST-2107253. The computations in this paper were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This research has made extensive use of NASA’s Astrophysics Data System Bibliographic Services.
Appendix A All-Sky Velocity Maps vs Distance

### v2

@@ 4.4 Local Effects of the LMC
To more closely examine the local velocity perturbations due to the LMC, we compare selected observed velocity trends to the simulations of GC19 in Figure 13. The observed datapoints in Figure 13 are directly measured from our stellar dataset, i.e., they are independent of the $v_{\mathrm{travel}}$ model described above. An alternative and valuable approach would be to examine the residuals of the data compared to the $v_{\mathrm{travel}}$ model, but we elect to show directly observed quantities here to highlight the simplest observable signatures that future studies should search for. The top panel shows the mean Galactocentric radial velocity along the $L_{\mathrm{MS}}$ coordinate, which is roughly aligned with the past orbit of the LMC (Nidever et al., 2008). The extent of this plot only contains stars in the southern Galactic hemisphere. The overall trend of decreasing $v_{\mathrm{GSR}}$ versus $L_{\mathrm{MS}}$ is a signature of the global reflex response (red line), but there are higher-order ‘wiggles’ in the trend from N-body MW+LMC simulations (blue lines) that signal local effects from the LMC. It is clear that even along the relatively precisely-measured radial velocity component, a much larger sample of stars will be required to disentangle the kinematic signature of the local wake from the global reflex motion.
@@ 5.1 Context and Limitations
Before discussing the conclusions of our work, it is important to highlight the limitations of our dataset and methodology. As Figure 2 illustrates, our MagE survey has expanded spectroscopic coverage of the outer halo to the southern hemisphere, a major improvement over past samples. However, the bottom panel of Figure 2 also shows that the MagE survey is much more efficient at identifying stars beyond $\gtrsim 60$ kpc than corresponding northern surveys (H3 and SEGUE). Therefore, there is a strong need to observe more $\gtrsim 60$ kpc stars in the northern hemisphere, to achieve homogeneous sky coverage at these distances. Regardless, we have demonstrated (see Figure 4 and Appendix D) that the deeper southern coverage should not negatively affect the main measurements in our paper.
@@
The $v_{\mathrm{travel}}$ measured here differs in the sign of $v_{\mathrm{Y}}$ compared to the past measurements of PP21 and YPP24. In Appendix A, we investigate this discrepancy in detail. Briefly, we find that the PP21 and YPP24 measurements appear to be significantly influenced by the presence of Sagittarius Stream (Sgr) stars in their datasets (as first noticed by A. Byström, private communication). These Sgr stars have an on-sky velocity distribution that closely mimics a dipole pointing in the quadrant opposing the LMC. In Appendix A, we demonstrate that after removing this Sgr contamination, the PP21 dataset does not have a detectable $v_{\mathrm{travel}}$ signature (Figure 15). Therefore, the most plausible explanation for the difference between our measurements and those from PP21 and YPP24 is the presence of Sgr contamination in the earlier datasets.
Another possible reason our measurement could differ from previous ones is the larger distance and sky coverage of our dataset. In Appendix B we test the effect of sky coverage by fitting limited subsamples of our data, finding that our SEGUE+H3+MagE dataset can recover $v_{\mathrm{travel}}$ even if only stars with $\delta>-20^{\circ}$ — observable from the northern hemisphere — are fitted, although the uncertainties are larger. However, if we only use SEGUE data, $v_{\mathrm{travel}}$ becomes undetected. Therefore we concur with YPP24 that $v_{\mathrm{travel}}$ can in principle be measured using only data observed from the northern hemisphere, but conclude that the increased sample size (and larger distance coverage) of the H3 and MagE datasets are vital contributors to our measurement.
@@
Our agreement with simulations is by no means perfect, with the amplitude of our measured $v_{\mathrm{Y}}$ being larger than the simulations of GC19 and the Sgr measurements of Vasiliev et al. (2021). One caveat is that our inclusion of southern data could make our measurements more susceptible to local halo perturbations of the LMC. Although the simulations of GC19 suggest that the all-sky velocity structure is quite smooth and dominated by the reflex motion itself, any deviations from the simulation (or unresolved substructure from the LMC halo itself) could bias our measurement. The LMC’s local influence, including the presence of stripped stars from the Clouds, can be explicitly modeled in the future.
@@
Finally, an implicit assumption in our measurements — and the comparison to simulations — is that there is a relatively smooth component of the MW halo, relative to which a reflex velocity can be measured. The $v_{\mathrm{travel}}$ measurement uncertainties quoted in this work are purely statistical, assuming a particular kinematic model. We have carefully removed known substructures including dwarf galaxies and the Sagittarius Stream from our dataset. However, additional unknown substructures may be present in the data. When fitting our $v_{\mathrm{travel}}$ model, we iteratively clip stars that disagree with the model by $>3\sigma$. These cuts should prevent small-scale undetected substructures from biasing our overall measurement. Regardless, given the emerging picture of the various constituents of the outer halo, it would be desirable for future simulations of the MW+LMC interaction to incorporate more realistic and structured models of the MW halo. In addition to including the Sagittarius Stream (e.g., Vasiliev et al., 2021), simulations could utilize the latest observations of the Galactic halo’s spatially-varying density profile (e.g., Han et al., 2022; Amarante et al., 2024) and velocity ellipsoid (e.g., Bird et al., 2021, Figures 11-12 in this work).
@@ 5.2 The Dynamical Disequilibrium of the Outer Halo
Prior to any detailed modeling, §$\S$3 and Figures 3-5 in this work clearly illustrate the dramatic departure from any notion of ‘dynamical equilibrium’ beyond $\approx 50$ kpc. In §$\S$4 we use a physical model to demonstrate that the kinematic disequilibrium of the Galactic outskirts can be explained — to first order — by the gravitational influence of the LMC. The kinematic influence of the LMC on our sample of stars is notably strong, suggesting an LMC dark matter halo on the massive side of recent estimates, $\gtrsim 15\%$ the mass of the Milky Way.
@@ 5.3 The Elusive Dark Matter Wake of the LMC
The dynamical friction wake of the LMC is predicted to leave imprints in both the density and kinematics of halo stars along the past orbit of the LMC (GC19; Belokurov et al., 2019; Conroy et al., 2021; Cavieres et al., 2024). Although this wake has been detected in the density of stars — forming the so-called ‘Pisces Plume’ overdensity — a confident detection of the wake’s kinematic signature remains elusive. In Figure 13 we search for the LMC’s local kinematic perturbation in radial and tangential velocities, over and above the bulk reflex motion. However, even in the relatively well-measured radial velocity direction, the detection is weak to non-existent given current observational errors.
@@ 6 Conclusions
We have assembled the largest all-sky dataset of $851$ luminous red giant stars from $40-160$ kpc with full 6D kinematics, including $\approx 300$ new stars in the southern hemisphere from our own tailored spectroscopic survey. In this work, we utilize this dataset to measure the motion of the inner galaxy relative to the outer halo, and find that it is induced by a massive and recently-infalling LMC. Our main conclusions are summarized below.
  • We construct an all-sky dataset of $851$ red giant stars beyond $40$ kpc using spectroscopy from H3, SEGUE, and our own MagE outer halo survey. All spectra are homogeneously analyzed with the same stellar parameter pipeline, resulting in the largest all-sky catalog of distant halo stars with 6D kinematic information (Figures 1-2).
@@
  • We fit the 6D positions and velocities of 821 stars from $12040<r_{\mathrm{gal}}<$ 120 to infer the amplitude and direction of the inner MW’s motion relative to distant halo stars. The apex direction of this travel velocity points towards the past orbit of the LMC, close to its present-day location (Figure 6, top). The amplitude of the travel velocity grows from $\sim 0$ at $r_{\mathrm{gal}}=40$ up to $50\pm 20$ km s-1 at $r_{\mathrm{gal}}=120$ kpc (Figure 6, bottom). A 3-D schematic of the measurement is shown in Figure 10 (animated version here).
  • The $v_{\mathrm{travel}}$ direction and amplitude measured in this work align with predictions from N-body simulations of the MW halo’s response to the LMC. The measured amplitude of the reflex velocity is most consistent with an LMC that is $\approx 15\%$ the mass of the Milky Way, based on the simulations of GC19 (Figure 6).
@@
  • The intrinsic velocity dispersion of the MW halo declines as a function of distance in along all three spherical coordinates. After accounting for the LMC-induced reflex motion, outer halo stars exhibit significantly radially-biased motions, with a rising profile of the velocity anisotropy parameter $\beta\approx 0.6-0.8$ from to $40-120$ kpc (Figure 12).
@@
We thank the referee, Michael Petersen, for thorough feedback that significantly improved this work. VC gratefully acknowledges a Peirce Fellowship from Harvard University. RPN acknowledges support for this work provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. CL acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 852839). It is a pleasure to thank Amanda Byström, Nitya Kallivayalil, David Nidever, Theo O’Neill, Ekta Patel, Michael Petersen, Yanjun Sheng, Eugene Vasiliev, and Rashid Yaaqib for insightful conversations and feedback. This work benefited from discussions at the 2024 “Milky Clouds over Manhattan” Workshop at the Flatiron Institute Center for Computational Astrophysics. We are indebted to the CfA and MIT telescope time allocation committees for enabling our long-term survey of the Galactic outskirts. We thank the staff at Las Campanas Observatory — including Yuri Beletsky, Carla Fuentes, Jorge Araya, Hugo Rivera, Alberto Pastén, Mauricio Martínez, Roger Leiton, Matías Díaz, Carlos Contreras, and Gabriel Prieto — for their invaluable assistance. The H3 Survey is funded in part by NSF grant NSF AST-2107253. The computations in this paper were run on the FASRC Cannon cluster supported by the FAS Division of Science Research Computing Group at Harvard University. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This research has made extensive use of NASA’s Astrophysics Data System Bibliographic Services.
Appendix A The PP21 Dataset
A key conclusion from our work is that the apex direction of the MW’s motion points towards a very recent location along the past orbit of the LMC (Figure 6). Although this apex direction agrees well with the results from Byström et al. (2024), it is somewhat different from the direction reported in PP21 and YPP24. In particular, the latter two works report an apex direction that points towards the opposing Galactic quadrant as the LMC (Figure 8). In this section, we examine the dataset used by PP21 to investigate the source of this difference. We use the K giant dataset from PP21 — although that work also fitted BHB stars and satellite galaxies, the K giants were by far the most constraining due to their low proper motion uncertainties. We find that the most plausible reason for our differing measurements is the presence of Sagittarius Stream stars in the PP21 and YPP24 datasets.
Figure 14: Comparing different angular momentum selections to identify and remove debris from the Sagittarius Stream. The entire sample from PP21 is shown in black. Stars that they identify as Sgr members using a spherical angular momentum selection are shown in orange, and stars that our linear angular momentum cut would identify as Sgr debris are shown in red. For reference, the Sgr simulation from Vasiliev et al. (2021) is shown in green.
PP21 and YPP24 utilized the angular momentum-based cuts from Peñarrubia & Petersen (2021) to identify Sgr stars, selecting stars within 3000 kpc km s-1 of the Sgr progenitor’s angular momentum: $[L_{x}=605,L_{y}=-4515,L_{z}=-1267]$ kpc km s-1. Figure 14 shows the K giant dataset from Petersen & Peñarrubia (2021), with stars satisfying their Sgr selection highlighted in orange. We only show stars within $20^{\circ}$ of the Sgr orbital plane, as defined by the Law & Majewski (2010b) coordinate frame. In red we show stars from their dataset that would fail our broader angular momentum selection drawn from Johnson et al. (2021), which removes all stars with $L_{\mathrm{Y}}<-2.5-0.3\times L_{\mathrm{Z}}$, where the angular momentum is in units of $10^{3}$ kpc km s-1. In the background, green points show Sgr stars from the simulations of Vasiliev et al. (2021), which were carefully fit to observations of the stream.
Figure 15: Illustrating the presence and impact of Sgr debris in the sample from PP21. Top: Galactocentric radial velocity as a function of Galactic longitude. Middle: On-sky distribution in Galactic coordinates. We show the clean sample from PP21 as black circles, highlighting with red crosses stars that our cuts would identify as Sgr debris. For reference, the Sgr simulation from Vasiliev et al. (2021) is shown (green). Bottom: Stars from PP21 that we would classify as Sgr debris, colored by Galactocentric radial velocity. The Vasiliev et al. (2021) simulation of Sgr stars is also shown with the same color scale.
Figure 14 shows that there is a marked overdensity of Sgr-like stars left in the PP21 dataset after their Sgr cuts are applied. Note that the observed stars span a wider swath of angular momentum space than the Vasiliev et al. (2021) simulations due to observational uncertainties, chiefly the distance uncertainty. To further illustrate this point, we show the clean PP21 dataset in Figure 15 after their Sgr cut has been applied. The top panel shows radial velocities as a function of Galactic longitude, and the middle panel shows an all-sky map in Galactic coordinates. In both panels, it is clear that a significant fraction of the K giant dataset fit by PP21 is coincident with Sgr debris. In particular, of the 548 stars in the clean PP21 dataset, our angular momentum cuts would have removed a further 165 stars as Sgr contaminants, 30% of the dataset fit by PP21.
The bottom panel of Figure 15 illustrates how the 30% Sgr contamination could bias the reflex motion apex direction inferred by PP21 (and by extension, YPP24). We show the 165 stars in their final sample that fall into the Sgr selection used in this work, colored by Galactocentric radial velocity. In the background are Sgr particles from the simulations of Vasiliev et al. (2021). The Sgr debris have a strong signature in radial velocities, which broadly appears like a dipole on the sky: moderately positive in the northern hemisphere, and strongly negative in the southern hemisphere. Furthermore, this dipole is oriented along the Sgr stream, with an apex direction (region of maximum negative Galactocentric radial velocity) in the Galactic quadrant opposing the LMC. This is the same quadrant as the apex direction inferred by PP21 (indicated in Figure 15, see also Figure 8). Clearly, the presence of these Sgr stars could significantly influence the apex direction inferred by PP21.
Figure 16: Posterior distribution of $v_{\mathrm{travel}}$ in Galactocentric coordinates inferred by fitting the PP21 dataset, before (top panel) and after (bottom panel) applying our angular momentum cuts to remove the $\approx 30\%$ residual Sgr contamination in that dataset. The contours show our measurements, and the diamonds show the reported measurements in PP21. Blue contours/markers correspond to $v_{\mathrm{travel,x}}$ on the x-axis and $v_{\mathrm{travel,y}}$ on the y-axis, and orange ones correspond the same in the X--Z plane.
To further demonstrate the influence of Sgr contamination, we fit the stars in the PP21 dataset with our $v_{\mathrm{travel}}$ model, and also fit their dataset after applying our more broad angular momentum selection (Figure 14). Following PP21, we use the distance-independent $v_{\mathrm{travel}}$ model here, fitting all stars in the dataset with a single set of $v_{\mathrm{travel}}$ parameters (see Appendix E for the analogous measurement on our own dataset). The resulting posterior distributions of $v_{\mathrm{travel}}$ are shown in Figure 16. The top panel shows fits to the PP21 dataset after applying their Sgr cuts, and the bottom panel shows the same after our Sgr cuts are applied to their dataset. Once our Sgr cuts are applied, the apparent $v_{\mathrm{travel}}$ signature in the PP21 dataset goes away, with $v_{\mathrm{travel}}$ being within $1\sigma$ of zero in all components. The apparent $v_{\mathrm{travel}}$ measured by PP21 (and by extension YPP24) therefore appears to be primarily driven by the kinematics of Sgr contaminants in those datasets, although some residual $v_{\mathrm{travel}}$ signal may remain.
The top panel of Figure 16 also shows that any differences with PP21 are not due to a mis-specification of our $v_{\mathrm{travel}}$ model, since we recover their quoted parameters within $1\sigma$ — the small disagreement in $v_{\mathrm{travel,x}}$ is likely due to subtle differences in the sample selection or sigma-clipping methodology. We conclude that our modeling framework is consistent with that of PP21 and YPP24, and the differences in our results are primarily due to Sgr contamination in the earlier datasets (see also Byström et al. 2024, who arrive at the same conclusion).
Appendix B The Importance of All-Sky Coverage
In Appendix A we argue that the primary reason our results differ from past measurements made by PP21 and YPP24 is our different strategies to remove contamination from the Sagittarius Stream, with apparent Sgr contamination visible in the earlier datasets (see Figure 15). However, our dataset differs from past data in two other ways: the inclusion of the H3 Survey adds a significant number of outer halo stars in the northern hemisphere, and the inclusion of our MagE survey fills in a missing southern Galactic quadrant (see Figure 2). In this section, we test the effect of sky coverage on our $v_{\mathrm{travel}}$ measurement.
We repeat the $v_{\mathrm{travel}}$ measurement described in §$\S$4 (incorporating the linear continuity model from $40-120$ kpc) on two subsets of our data. The first subset emulates the selection function of surveys observed from the northern hemisphere (e.g., H3 and SEGUE): $\delta>-20^{\circ}$. The resulting subset is predominantly composed of H3 and SEGUE stars, although some MagE stars at equatorial latitudes are also present. The second subset emulates the SEGUE-only dataset used by PP21, albeit with our updated MINESweeper stellar parameters and broad cuts to remove Sagittarius contamination.
Figure 17: Posterior distribution of $v_{\mathrm{travel}}$ in Galactocentric coordinates, inferred by fitting limited subsamples of our dataset. We show the measurement from using our full dataset (solid contour, identical to Figure 9), from using only stars observable from the northern hemisphere (dashed), and from using only data from the SEGUE survey (dotted).
The resulting measurements of $v_{\mathrm{travel}}$ in Galactocentic coordinates are shown in Figure 17, evaluated at $r_{\mathrm{gal}}=100$ kpc analogous to Figure 9. Using north-only data — including our SEGUE, H3 and MagE data — produces results somewhat consistent with our full dataset, albeit with larger uncertainties. Furthermore, the $v_{\mathrm{travel,x}}$ component is mildly biased relative to the full fit, and the $v_{\mathrm{travel,y}}$ signature is slightly weaker.
On the other hand, using SEGUE-only data — after appropriately removing Sagittarius contamination using our angular momentum cuts (see Figure 14) — results in a non-detection of the $v_{\mathrm{travel}}$ signature. Note that this is analogous to the measurement shown in Figure 16 in that it only uses SEGUE data, but with some differences. We use our own MINESweeper stellar parameters rather than the Xue et al. (2015) parameters used by PP21. Furthermore, in Figure 16 we fit a single $v_{\mathrm{travel}}$ at all distances for consistency with PP21, whereas in Figure 17 we use the full linear continuity model. There are less stars in our SEGUE dataset than in PP21 since we apply more stringent SNR cuts (SNR $>10$ per pixel) to ensure reliable stellar parameters. We have verified that there are no major systematic differences between our radial velocities and distances compared to the parameters from Xue et al. (2015). Despite these differences, the conclusion from both these tests is similar: after Sagittarius contamination is removed, there does not appear to be a strong $v_{\mathrm{travel}}$ signature in the SEGUE-only dataset.
In summary, the tests shown in Figure 17 demonstrate that although it is feasible to detect $v_{\mathrm{travel}}$ — albeit with mildly biased orientations — using only data observed from the northern hemisphere (as argued by YPP24, and recently shown by Byström et al. 2024 using DESI data), our SEGUE dataset alone appears insufficient to reliably measure $v_{\mathrm{travel}}$. The inclusion of H3 and MagE data substantially improves the north-only measurement, and the best measurement (unsurprisingly) comes from our all-sky dataset that incorporates all three surveys. These tests also demonstrate that the reflex response is a sensitive and challenging measurement that can be affected by a variety of factors including sky coverage, distance coverage, and contamination by substructure.
Appendix C Fitting Data-Matched MW+LMC Simulation
When fitting any model to data, it is important to consider the influence of the data’s selection function. Our dataset has the advantage of sampling the entire sky, an improvement over past measurements. Furthermore, our MagE dataset systematically spans larger distances than past halo datasets. However, as the top panel in Figure 2 illustrates, this produces an on-sky asymmetry in the typical distance towards stars in our dataset. Note that for our $v_{\mathrm{travel}}$ measurement, we restrict the data to $40-120$ kpc, which somewhat mitigates this asymmetry. However, it is important to verify that this selection effect does not affect the conclusions of this work.
Throughout this work, we have shown for comparison $v_{\mathrm{travel}}$ measurements made by fitting the simulations of GC19 (e.g., see Figure 6). For this Appendix, we sub-sample the simulation particles to match the spatial distribution and size of our final dataset. Specifically, for each of the $821$ stars in our final observed sample, we select the nearest particle (in 3D Galactocentric coordinates) from the GC19 LMC3 simulation, which has a $1.8\times 10^{11}\,M_{\odot}$ LMC. This produces a mock dataset that exactly matches the spatial selection function of our observed dataset. Since distances are expected to be our dominant source of measurement uncertainty, we assume 10% distance uncertainties on this mock dataset.
[…]
We fit the data-matched simulation using the $v_{\mathrm{travel}}$ model described in §$\S$4, and show the resulting posterior contours in Figure 18. The $v_{\mathrm{travel}}$ inferred from fitting the simulation without any imposed selection function is also shown. We also verify that the distance-dependence of $v_{\mathrm{travel}}$ is well-captured, following the intrinsic behavior of the simulation well within $1\sigma$ uncertainties. Finally, we verify that our result is not affected by our broad angular momentum cut to remove Sgr contamination (based on Johnson et al. 2020). Although Johnson et al. (2020) applied this cut to stars across the sky, we only apply it to stars within $25^{\circ}$ of the Sgr orbital plane, and removes $\approx 15\%$ of field stars from the simulated dataset. However, removing these stars does not meaningfully alter the best-fit $v_{\mathrm{travel}}$ parameters. These tests support our claim that the size and spatial selection function of our dataset — and our strategy to remove Sgr contamination — does not limit or bias its ability to measure $v_{\mathrm{travel}}$.
For ease of comparison to future works, we include in Table 3 our best-fit parameters to the LMC3 simulation from Garavito-Camargo et al. (2019). To determine these parameters, we randomly sub-sampled $10,000$ particles from the simulation, with no selection function or uncertainties applied. These parameters (and their analogs for the other three simulations from GC19) are used as the reference comparisons throughout this work.
Table 3: Summary of the best-fit parameters from applying our $v_{\mathrm{travel}}$ continuity model to a random sub-sample of $10,000$ stars from the LMC3 simulation of GC19.
Parameter $r_{\mathrm{gal}}=40$ kpc $r_{\mathrm{gal}}=120$ kpc
$v_{\mathrm{travel}}$ [km s-1] $11$ $51$
$\ell_{\mathrm{MS,apex}}$ [deg] $-25$ $-48$
$b_{\mathrm{MS,apex}}$ [deg] $8$ $4$
$\langle v_{\mathrm{r}}\rangle$ [km s-1] $-5$ $-20$
$\langle v_{\mathrm{\phi}}\rangle$ [km s-1] $-1$ $-3$
$\langle v_{\mathrm{\theta}}\rangle$ [km s-1] $8$ $9$
${\sigma_{\mathrm{v,r}}}$ [km s-1] $115$ $87$
${\sigma_{\mathrm{v,\ell}}}$ [km s-1] $90$ $62$
${\sigma_{\mathrm{v,b}}}$ [km s-1] $85$ $66$
Appendix D All-Sky Velocity Maps vs Distance
Figure 19: All-sky maps of the mean Galactocentric radial velocity $v_{\mathrm{GSR}}$ using our dataset, smoothed on $45^{\circ}$ scales. These figures are computed identically to Figure 3, but are split into four disjoint $r_{\mathrm{gal}}$ bins. On-sky positions of stars in each distance bin are shown as black circles.